1. What is the magnitude of the electric field at a field point 2 m from a point charge q0 = 4 nC?  

 


2.  Determine the electric field of a –60-nC charge that is experiencing an electric force of 2.75 x 10–4 N.   

 

3. Suppose two point charges, q1 = + 43 nC and q2 = –17 nC, are separated by a distance of 6 cm. Determine the:

(a) electric force interacting between the two charges; and

(b) electric field of each charge.


4. Determine the electric field strength and potential in air at a distance of 3 cm from a charge of 5 X 10-8 C.


1
Expert's answer
2022-02-15T10:58:24-0500

1)


"\begin{vmatrix}
 E \\
 
\end{vmatrix}=\frac{KQ}{r^2}"


"=\frac{9×10^9×4×10^{-9}}{2^2}=9N/C"


2)


"\begin{vmatrix}
 E \\
 
\end{vmatrix}" "=\frac{F}{q}"


"=\frac{2.75×10^{-4}}{60×10^{-9}}=4583.3N/C"


3)

a)

"\begin{vmatrix}
 F \\
 
\end{vmatrix}=\frac{Kq_1q_2}{r^2}"


"=\frac{9×10^9×43×17×(10^{-9})^2}{(6×10^{-2})^2}"

"=1.8275\u00d710^{-3}N"


b)

"\begin{vmatrix}
 E_{q_1} \\
 
\end{vmatrix}=\frac{F}{q}=\frac{1.8275×10^{-3}}{17×10^{-9}}"


"=1.075\u00d710^5N\/C"


"\begin{vmatrix}
 E_{q_2} \\
 
\end{vmatrix}=\frac{1.8275×10^{-3}}{43×10^{-9}}=4.25×10^4N/C"



4)


"E=\frac{KQ}{r^2}=\frac{9×10^9×5×10^{-8}}{(3×10^{-2})^2}=5×10^5N/C"


"V=\frac{KQ}{r}=\frac{9×10^9×5×10^{-8}}{(3×10^{-2})}=1.5×10^4V"